Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2007-12-03
ISS On-Orbit Status 12/03/07

Underway: Week 7 of Increment 16.

FE-1 Malenchenko continued his support of his first experiment session with the Russian TEKh-20 Plasma Crystal-3+ (Plazmennyi-Kristall/PK-3+) payload by activating the turbopump in the Service Module (SM)'s Transfer Compartment (PkhO) for keeping the vacuum chamber (ZB) in the SM Work Compartment (RO) evacuated. The turbopump will be deactivated tonight at ~4:25pm EST. (Main objective of PK-3 is to study dust plasma wave propagation and dispersion ratio at a specified power of HF discharge, pressure, and a varied number of particles.)

CDR Whitson and FE-2 Tani assisted the ground in deactivating the CDRA (Carbon Dioxide Removal Assembly) at 2:25am-3:25am EST, used to support yesterday's SPHERES experiment. With cooling no longer required, the crew also disconnected the LTL (Low Temperature Loop) jumper of the ITCS (Internal Thermal Control System) from the CDRA rack.

At ~3:05am, in support of HDTV (High Definition TV) playback & downlink by ground commanding, FE-2 activated the Photo/TV MPC (Multi-Purpose Converter). (MPC was turned off again by Tani at ~1:20pm.)

Dan then also set up the Ku-band for downlinking analog video signals (streaming video) via U.S. OCA from the Russian TVS television system in a test of the relatively new KL-211 MPEG-2 TV Encoder in the new JSL (Joint Station LAN).

The CDR prepared the CSLM-2 (Coarsening in Solid-Liquid Mixtures-2) experiment for operation, turned on the SAMS ICU (Space Acceleration Measurement System/Interim Control Unit) in LAP2 Rack, and started CSLM-2 by activating heating of the sample.

Peggy and Dan spent time in the Joint Airlock to continue EVA preparations, today verifying and configuring tools for the SARJ (Solar Alpha Rotary Joint) inspection on 1E EVA-2, then performing the regular checkout and test of the SAFER (Simplified Aid for EVA Rescue) units to be used for the spacewalks by Walheim, Schlegel and Love (#1006 & #1007).

Later, the crew conducted a joint review of uplinked STS-122/1E spacewalk procedures and will link up with EVA personnel at MCC-Houston at ~4:00pm via Ku- & S-band to discuss details.

Peggy and Dan each had ~90 min set aside to prepack cargo items for transfer to STS-122/Atlantis for return to Earth. The FE-2 used an additional 15 min for prepacking the failed IVCPDS (Intravehicular Charged Particle Directional Spectrometer) hardware for return. (The IVCPDS, found to be hard failed on 10/30, was a secondary radiation detection measurement tool; the primary radiation measurement tool is the TEPC (Tissue Equivalent Proportional Counter).)

The CDR conducted the microbial (bacterial & fungal) 'T+5 Day' analysis of air samples collected on 11/28 by Dan Tani with the MAS (Microbial Air Sampler) kit in Lab, Node and SM. (The sampling analysis is performed once per month for the first three months, and once every three months thereafter. Bacterial and fungal air samples are taken at two locations in each module. The colony growth on the sampling slides is analyzed after five days of incubation in four Petri dishes. For onboard visual analysis of media slides from SSK (Surface Sampling Kit), MCDs (microbial capture devices) from WMK (Water Monitoring Kit), coliform detection bags from MWAK (Microbial Water Analysis Kit), and Petri dishes from MAS, the crew has a procedure for visual inspection of samples for bacterial and fungal colony growths after appropriate incubation periods.)

Continuing the current round of monthly preventive maintenance of RS (Russian Segment) ventilation systems, FE-1 Malenchenko inspected and cleaned the 'Group B' (B1 & B2) fan grills in the SM, followed later by reporting to TsUP-Moscow on equipment status and any presence of moisture in work areas.

Yuri also performed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.)

Dan handled the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Malenchenko set up the "Urolux" equipment, hooked it up to the Power Supply Center (BP Tsentr) and ran the periodic functional health test and calibration on it, then deactivating the hardware.

Later, the FE-1 unstowed and installed the equipment for the periodic Russian PZE-MO-10 "Hematokrit" testing that is scheduled tomorrow for him. (MO-10 measures the hematocrit (red blood cell mass) value of the blood (it is a well-known phenomenon of space flight that red blood cell mass {normal range: 30-45%} tends to go down over time).)

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR), TVIS treadmill (FE-1, FE-2), RED resistive exerciser (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Dan Tani copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~1:05pm, Peggy and Yuri had a 15-min teleconference with ground specialists to discuss the images downlinked from their recent (11/29) Shuttle RPM (R-bar Pitch Maneuver) skill training. (The skill training prepares the crew for the bottom side mapping of the Orbiter at the arrival of STS-122/1E. During the RPM at ~600 ft from the station, the ISS crew will have only ~90 seconds for taking high-resolution digital photographs of all tile areas and door seals on the Orbiter, to be downlinked for launch debris assessment. Thus, time available for the shooting will be very limited, requiring great coordination between the two headset-equipped photographers and the Shuttle.)

At ~9:35am, Whitson and Tani participated in two live PAO TV interviews with media in Chicago, one with WGN Radio (Spike O'Dell), the other with WJMK Radio (Steve Dahl).

At 2:30pm, Dan is scheduled for a CDC (Crew Discretionary Conference).

Lab/Node-2 Vestibule Leak Check Update: The second Fine Leak check on the Lab/Node-2 vestibule conducted by Peggy Whitson on 12/1 has passed. The indicated pressure drop of ~9 mmHg in 7 hours is equivalent to a loss of 0.13 pound-mass of air at 14.7 psia. This is far below the limit criterion used, resulting in the conclusion that there is no leak. The vestibule was completely repressurized by equalization with the ISS stack and is ready for 1E.

SSRMS Pre-launch Checkout Update: The regular pre-launch checkout of the Canadian SSRMS (Space Station Remote Manipulator System) went without issue. The robotarm is now properly positioned for 1E, having 'walked off' the Lab PDGF (Power & Data Grapple Fixture) and changed its base to the Node-2 PDGF.

No CEO (Crew Earth Observation) photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use